Given the vectors: 𝑟 = 〈9, −7, 7〉, 𝑣 = 〈3, −6, 9〉, 𝑤 = 〈−3, −6, 6〉. Find (𝑟 + 𝑣) ∙ 𝑤
Answer by Edwin McCravy(20054)

(𝑟 + 𝑣) ∙ 𝑤 =

(〈9, −7, 7〉 + 〈3, −6, 9〉) ∙ 〈−3, −6, 6〉 =

〈9+3, −7-6, 7+9〉 ∙ 〈−3, −6, 6〉 =

〈12, −13, 16〉 ∙ 〈−3, −6, 6〉 =

(12)(-3) + (-13)(-6) + (16)(6) =

-36 + 78 + 96 =

138

Edwin